OpenStudy (anonymous):

Conduct an experiment of throwing two dice. Find the probability that the sum of the two dice is 2. a. 1/36 = 0.03 b. 1/18 = 0.06 c. 1/12 = 0.08 d. 1/9 = 0.11

undeadknight26 (undeadknight26):

hmm im guessing this is a 6 sided dice?

OpenStudy (anonymous):

I think so! @undeadknight26

undeadknight26 (undeadknight26):

hmm sadly i must go rest...i just found out like 2 seconds ago... @agent0smith @AccessDenied @whpalmer4 @Destinymasha guys can help?

OpenStudy (anonymous):

1/6 *1/6 = 1/36

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Why is it that the answer? how did you made it? @jN7

OpenStudy (anonymous):

you need to roll a 1 on both dice to get 2, the probability of getting a 1 on a roll is 1/6, therefore 1/6 * 1/6 =36

OpenStudy (anonymous):

sorry typo, 1/6*1/6=1/36

OpenStudy (accessdenied):

If the problem were more complicated, such as rolling a sum divisible by 3, I would recommend a table. The table would represent all possible combinations you can roll. Label the columns 1 through 6 and the rows 1 through 6. The box corresponding to one row and one column you would write the sum of the row value and column value. The total boxes would be 6x6 = 36. You then highlight the boxes that support the probability (in this case, one box row 1 col 1 has entry 2). Take the division: 1 (number of wanted outcomes) divided by 36 (total outcomes). 6 5 4 3 4 ... 2 3 4 ... 1 (2) 3 4 ... 1 2 3 4 5 6 Again, it is extraneous for this problem although it can be more useful for different types of questions. "Sum greater than 6, etc."

OpenStudy (whpalmer4):

Here are all the possible rolls of 2 6-sided dice: \[\begin{array}{cccccc} \{1,1\} & \{1,2\} & \{1,3\} & \{1,4\} & \{1,5\} & \{1,6\} \\ \{2,1\} & \{2,2\} & \{2,3\} & \{2,4\} & \{2,5\} & \{2,6\} \\ \{3,1\} & \{3,2\} & \{3,3\} & \{3,4\} & \{3,5\} & \{3,6\} \\ \{4,1\} & \{4,2\} & \{4,3\} & \{4,4\} & \{4,5\} & \{4,6\} \\ \{5,1\} & \{5,2\} & \{5,3\} & \{5,4\} & \{5,5\} & \{5,6\} \\ \{6,1\} & \{6,2\} & \{6,3\} & \{6,4\} & \{6,5\} & \{6,6\} \\ \end{array}\] If you count them, you'll see that there are 36 possible rolls, and only 1 of them produces a sum of 2 (affectionately known as "snake eyes" by the craps players among us) so the probability is 1/36.